What equation has a slope of 2 and contains the point (4,-4)?
Answer by texttutoring(324)

Use y=mx+b with m=2, x=4, and y=-4. Solve for b:

-4 = 2(4) +b
-4 = 8 + b
-12 = b

y=2x - 12 is the equation.